A crocodile lays 35 eggs. Each egg has a probability of 0.8 of surving what are the probabilities of all eggs surving? None of the eggs surving? Exactly 33 of the eggs surviving, 33 or more eggs surviving? Less than 33 egs surviving
1
Expert's answer
2011-05-19T07:36:08-0400
P(all eggs surviving)= 0.835
P(none of eggs surviving)= (1-0.8)35= 0.235
P(exactly 33 of eggs surviving)= C(35, 33)*0.833*0.22
P(exactly 33 or more of eggs surviving)=
= C(35, 33) 0.833*0.22+ C(35, 34) 0.834*0.2+ C(35, 35) 0.835= (35*34)/2*0.833*0.22+35*0.834*0.2+ 0.835
P(Less than 33 egs surviving)= 1- P(exactly 33 or more of eggs surviving)